### v1

# The disk precession in a Be star-magnetar binary and its application to the rotation measure of FRB 20201124A

Abstract
Fast radio bursts (FRBs) are bright, millisecond-duration radio bursts with poorly known origins. Most FRB sources are detected only once, while some are repeaters. Variation patterns observed in the rotation measure (RM) of some repeaters — indicate that the local magneto-ionic environments of these FRB sources are highly dynamic. It has been suggested that a Be star-magnetar binary system is a possible origin for such variation. FRB 20201124A is notable among these sources since it is the most active one and exhibits substantial temporal variations of RM measured by the Five-hundred-meter Aperture Spherical radio Telescope (FAST). The physics behind this long-term behavior is poorly understood. Here we propose that, within the framework of the Be star-magnetar binary scenario, the observed variation of RM is attributed to a combination of orbital motion and the precession of the circumstellar disk of the Be star. While a $\sim 785$-day precession of the disk contributes to the observed decrease in the amplitude of the variation, our model predicts that the amplitude oscillates with this period.
Fast radio bursts — Rotation measure — Binaries
show]leiwh@hust.edu.cn
show]zouyc@hust.edu.cn
show]fayinwang@nju.edu.cn
1 Introduction
FRBs are a class of cosmological radio transients with millisecond durations and high luminosity (Lorimer_2007Sci...318.. [^7] L; Thornton_2013Sci...341...53T; Wu2024). Although the origin of FRBs remains poorly understood, the detection of FRB 200428 from the Galactic magnetar SGR J1935+2154 reveals that magnetars are a viable source of FRBs (Bochenek_2020Natur.587...59B; CHIME_2020Natur.587...54C). Many researchers indicate that the FRB emission region is likely to be within the magnetosphere of the magnetar (Kumar_2017MNRAS.468.2726K; Yang2018; Nimmo_2025Natur.637...48N). Approximately 900 FRB sources have been detected to date, most of them are detected only once, whereas others exhibit repeated emissions (Spitler_2016Natur.531..202S; CHIME_2019ApJ...885L..24C; CHIME_2023ApJ...947...83C). Long-term observations of FRB repeaters allow for the investigation of their properties, e.g., dispersion measure (DM), rotation measure (RM), polarization, burst rate, and waiting time, helping to reveal the origin of FRBs.
Over a decade of observations of FRBs, several FRBs exhibit dramatic RM variations and oscillations (Hilmarsson_2021MNRAS.508.5354H; Xu_2022Natur.609..685X; Kumar_2023MNRAS.526.3652K; Anna-Thomas_2023Sci...380..599A; Mckinven_2023ApJ...950...12M; Ng_2025ApJ...982..154N). Since the RM is related to the magnetic field and plasma density, the temporal variation may indicate the evolution of the local magneto-ionic environment around the FRB source.
Among these FRBs, FRB 20201124A attracts significant attention. It is an extremely active repeater and exhibits peculiar variations in its RM. From the observations by FAST between April and June 2021 (Xu_2022Natur.609..685X), the RM of FRB 20201124A exhibits a pattern of rapid increase, sharp drop, and subsequent re-ascent over $\sim$ 40 days, with the maximum RM variation reaching $\sim$ 500 $\rm{rad/m^{2}}$. The variation stage is overlaid by a series of oscillations, and followed by a stable stage lasting for $\sim$ 20 days. After subtracting the RM averaged over $\sim$ two-month observations (from 1 April to 11 June, 2021), the variations show that the local RM contribution of FRB 20201124A changes its sign (Wang_2022NatCo..13.4382W).
After the 16-day periodically modulated activity for FRB 20180916B was found, the orbital motion of binary systems is proposed (Ioka_&_Zhang_2020ApJ...893L..26I; Katz_2020MNRAS.494L..64K; Sridhar_2021ApJ...917...13S; Pleunis_2021ApJ...911L...3P). However, these models cannot account for the frequency-dependent active window (2021Natur.596..505P). For the observed RM variation of FRB 20201124A, Wang_2022NatCo..13.4382W examined a Be star and magnetar binary scenario and suggested that the orbital motion causes a variation in the local magneto-ionic environment along the line of sight (LoS), resulting in the sign-changing RM. This model is supported by the following studies. Xu_2025arXiv250506006X claimed the detection of a $\sim$ 26-day periodicity with high significance in the RM observations of FRB 20201124A from 2021 to 2022. Du_2025arXiv250312013D detected a $\sim$ 1.7 s periodicity with high significance in the times of arrival (TOAs) of FRB 20201124A bursts observed on MJD 59310 and MJD 59347; such a period is explained by the spinning of the potential magnetar central engine of FRB 20201124A. Zhang_2025arXiv250517880Z further notes that the two days with detected TOAs periodicity bookend the RM variation phase of FRB 20201124A, and they might be the two observational windows when the magnetar central engine crossed the Be star disk. In addition to FRB 20201124A, Liang_2025arXiv250510463L reported the existence of a $\sim$ 200-day periodicity in the RM of FRB 20220529 from 2022 to 2025. The RM variation of FRB 20180916B may be caused by the stellar wind of a massive companion (Zhao2023). These potential periodicities in RM are likely to be related to the orbital motion of binary star systems (Rajwade2023; ZhangB2025; Wang2025).
Super-orbital brightness variations — periodic brightness oscillations far longer than the binary’s orbital period in multi-wavelength bands — have been observed in many Be star-neutron star binaries (Ogilvie_2001MNRAS.320..485O; McGowan_2008MNRAS.384..821M; Rajoelimanana_2011MNRAS.413.1600R) — specifically a Be/X-ray binary (Negueruela_1998A&A...336..251N; Coe_2005MNRAS.356..502C; Liu_2005A&A...442.1135L; Haberl_&_Sturm_2016A&A...586A..81H). While the driving mechanism of this periodic behavior is unclear, many studies indicate that it could be induced by Be disk precession (Rajoelimanana_2011MNRAS.413.1600R; Martin_&_Franchini_2021ApJ...922L..37M; Martin_2023MNRAS.523L..75M; Martin_&_Charles_2024MNRAS.528L..59M; Chen_2024ApJ...973..162C). Assuming that the rotation axis of the Be star is precessing with a superorbital period, and according to Martin_2023MNRAS.523L..75M, the decretion disk can remain locked to the equatorial plane of the Be star; the node of the disk plane and the neutron star orbit will precess with the star. Since the brightness variations of the Be star essentially reflect changes in the disk density along the LoS between the Be star and the observer, the disk precession can lead to the periodic modulation of the density profile along the line of sight, producing the super-orbital modulation in the observed brightness (Alcock_2001MNRAS.321..678A; Rajoelimanana_2011MNRAS.413.1600R; Martin_&_Charles_2024MNRAS.528L..59M). Moreover, X-ray bursts emitted when the neutron star interacts with the disk may also have their brightness modulated by the precession of the Be star and the disk (Alcock_2001MNRAS.321..678A). The magnitude of FRB 20201124A RM varies significantly, which cannot be explained by a simple binary orbital motion. The above studies motivate us to investigate the precession of the Be star disk in binary systems modulating the RM variations of FRBs.
In this work, we present a Be star-magnetar model with a precessing disk to elucidate the long-term RM variations in FRB 20201124A. The paper is organized as follows. The observations and data reduction of FRB 20201124A are presented in Section 2. The model of a binary system with a precessing disk is described in Section 3. The fitting results are shown in Section 4. In Section 5, we draw conclusions and discuss other FRB sources with similar RM variations.
2 FRB 20201124A
FRB 20201124A is a highly active repeating FRB source. By April 2021, FRB 20201124A entered an extremely active phase (Xu_2022Natur.609..685X), making it one of the most active known FRBs. In addition, FRB 20201124A is found to be associated with a persistent radio source (PRS), which supports the magnetar origin of this FRB (Bruni_2024Natur.632.1014B). Since 2021, FAST has been monitoring this source, covering the 1.0–1.5 GHz frequency range, and has detected thousands of bursts during the three active cycles (Xu_2022Natur.609..685X; Jiang_2022RAA....22l4003J; Niu_2022RAA....22l4004N; Xu_2025arXiv250506006X). The three FAST observation epochs are hereinafter referred to as E1, E2, and E3 for brevity. Only the bursts with high signal-to-noise ratio $S/N$ are adopted to measure the RMs by using a Q-U fitting method (Desvignes_2019Sci...365.1013D).
In E1, FAST monitored FRB 20201124A from April 1, 2021, to June 11, 2021 (UT). During a total observing time of 91 hours, 1,863 bursts were detected. A total of 1,103 bright bursts with a $S/N>30$ are adopted to measure the RM (Xu_2022Natur.609..685X). Intriguingly, the RM of the adopted bursts shows dramatic variation in a short term ($\sim 60$ days), and the maximum variation can reach about 500 $\rm{rad\,m^{-2}}$, making it the first FRB observed with rapid RM variation in such a short timescale (Xu_2022Natur.609..685X; Wang_2022NatCo..13.4382W).
E2 spans from 25 September 2021 to 17 October 2021 (UT), with a total observation of 19 hours. However, a total of 881 bursts was observed during the first four days, i.e., 25-28 September 2021, whereas no bursts were detected later in this epoch (Zhou_2022RAA....22l4001Z; Zhang_2022RAA....22l4002Z; Jiang_2022RAA....22l4003J; Niu_2022RAA....22l4004N). Among these 881 bursts, a total of 536 bursts with $S/N\geq 50$ are adopted to measure RM. The RM of the E2 bursts remains at the average level of the E1 bursts, but displays a feature of continuous drop. swinging and orthogonal jumping PAs are also found in some of these bursts (Niu_2024ApJ...972L..20N).
E3 is conducted by FAST from 2 February 2022 to 26 May 2022, during the third active cycle. In this epoch, 1,467 bright bursts with $S/N>20$ are detected in the total observing time of 34.3 hours (Xu_2025arXiv250506006X). The RM data of the E3 bursts also show a variation feature within several tens of days, but at a level of $\sim 100\,\rm{rad\,m^{-2}}$, which is a few times lower than that of epoch one. Xu_2025arXiv250506006X searched for the periodicity of the RM variation in the E3 bursts, and a period of $\sim 30$ days is found with high significance. We note that the E3 RM data have a lower background level compared to that of E1 and E2, which might be due to the long-term evolution of the magneto-ionic environment.
In this work, we use the FRB 20201124A RM data of E1, E2, and E3 (denoted as $\rm{RM_{E1}}$, $\rm{RM_{E2}}$, and $\rm{RM_{E3}}$ below, respectively). The data of $\rm{RM_{E1}}$ are provided by (Xu_2022Natur.609..685X), and the data of $\rm{RM_{E2}}$ and $\rm{RM_{E3}}$ are obtained from the plots in Xu_2025arXiv250506006X using the online data extraction tool WebPlotDigitizer (Marin_2017sf2a.conf..113M). To characterize RM variations, we perform some pre-processing operations on $\rm{RM_{E1}}$, $\rm{RM_{E2}}$, and $\rm{RM_{E3}}$. For each episode (episode $i$), we first bin $\rm{RM_{E\textit{i}}}$ by MJD and derive the mean RM value in each MJD bin (after this step, the data of episode $i$ are then denoted as $\rm{RM_{dE\textit{i}}}$). Secondly, we derive the mean for $\rm{RM_{dE\textit{i}}}$, which gives $\overline{\rm{RM}}_{\rm{dE1}}=-583.95\,\rm{rad}\,m^{-2}$, $\overline{\rm{RM}}_{\rm{dE2}}=-593.35\,\rm{rad}\,m^{-2}$, and $\overline{\rm{RM}}_{\rm{dE3}}=-666.45\,\rm{rad}\,m^{-2}$. Finally, we can obtain the RM variations $\Delta\rm{RM}_{E1}$, $\Delta\rm{RM}_{E2}$, and $\Delta\rm{RM}_{E3}$ by calculating $\Delta\rm{RM}_{E\textit{i}}=\rm{RM_{dE\textit{i}}}-\overline{\rm{RM}}_{dE\textit{i}}$.
3 Be star-magnetar binary with disk precession
3.1 Be star-magnetar binary
We consider a Be star-magnetar binary system. The stellar outflow from the Be star can be particularly significant in the equatorial region, where the dense flow usually forms a decretion disk. The density of the decretion disk decreases with increasing radius, and the base density $\rho_{0}$ typically ranges from about $10^{-12}$ to $10^{-10}\,\rm{g\,cm^{-3}}$ (Porter_1999A&A...348..512P; Gies_2007ApJ...654..527G; Rivinius_2013A&ARv..21...69R; Wang_2022NatCo..13.4382W).
Here, we consider an axisymmetric disk and describe the density profile as (Wang_2022NatCo..13.4382W)
$\rho(r,z)=\rho_{0}(\frac{r}{R_{\star}})^{-\beta}{\rm{exp}}[-(\frac{D}{H(r)})^{2}],$ (1)
where $r$, $D$, $R_{\star}$, and $\beta$ are the distances from the stellar center, the heights from the stellar equatorial plane, the stellar radius, and the density slope index, respectively. In the case of a Be star, $\beta$ usually ranges from $2\sim 4$. In addition, $H(r)$ is the vertical scale height of the disk, which is determined by (Bjorkman_1997LNP...497..239B; Rivinius_2013A&ARv..21...69R)
$H(r)\simeq c_{\rm s}(\frac{r}{GM_{\star}})^{1/2}r$, (2)
where $G$ is the gravitational constant, $M_{\star}$ is the stellar mass, and $c_{\rm s}$ is the sound speed in the disk. For an isothermal gas disk, $c_{\rm s}$ can be given by
$c_{\rm s}=\sqrt{\frac{kT}{\mu m_{\rm H}}},$ (3)
where $k$, $T$, $\mu$, and $m_{\rm H}$ are the Boltzmann constant, the isothermal temperature, the mean molecular weight, and the atomic mass of hydrogen, respectively.
Previous studies suggest that a magnetic field exists in the circumstellar disk (Reiners_&Mohanty_2012ApJ...746...43R; Pudritz_2014IAUS..302...10P; Tsukamoto_2016PASA...33...10T; Ohashi_2025NatAs...9..526O). Here, we adopt a simplified toroidal magnetic field (Johnston_1996MNRAS.279.1026J; Brauer_2017A&A...607A.104B)
$\textbf{B}(\textbf{r})={\rm{B_{0}}}(\frac{R_{\star}}{r})\vec{e}_{r}\times\vec{e}_{\rm{disk}},$ (4)
where $\rm{B_{0}}$, r, $\vec{e}_{r}$, and $\vec{e}_{\rm{disk}}$ are the magnetic field strengths at the stellar surface, the vector pointing from the stellar center to the field point, the unit vector of r, and the unit normal vector of the Be star disk, respectively.
The magnetar orbits the Be star along an elliptical path following Keplerian motion. The semi-major axis of the orbit can be expressed as
$a=\left[\frac{G(M_{\star}+M_{\rm{NS}})P_{\rm{orb}}^{2}}{4\pi^{2}}\right]^{1/3},$ (5)
[…]
We assume that the magnetar continuously emits FRB in its magnetosphere while moving along its orbit. The FRB propagating along the observer’s LoS may pass through the outflow material of the star, mainly the decretion disk. Thus, the density profile and magnetic field along the LoS in the disk determine the observed RM:
${\rm{RM}}=8.1\times 10^{5}\int_{\rm{LoS}}\frac{n_{\rm e}(l)\cdot\textbf{B}_{\parallel}(l)}{(1+z(l))^{2}}\,\textit{dl}\,\,\,\rm{rad\,m^{-2}},$ (6)
where $\textbf{B}_{\parallel}$ is the magnetic field component along the LoS (in units of Gauss), $z$ is the redshift, and $n_{\rm e}$ is the number density of free electrons (in units of $\rm{cm^{-3}}$). Assuming that the disk is mainly composed of hydrogen, the number density of the free electrons for the gas temperature $T$ is calculated from the Saha ionization equation:
$\frac{n_{\rm H^{+}}\cdot n_{\rm e}}{n_{\rm H}}=\frac{2g_{\rm H^{+}}}{g_{\rm H}}\left(\frac{2\pi m_{\rm e}kT}{h^{2}}\right)^{3/2}e^{-\chi_{0}/(kT)},$ (7)
where $g_{\rm H^{+}}$ and $g_{\rm H}$ are the statistical weights of the proton and neutral hydrogen atom, respectively (we set $g_{\rm H^{+}}=g_{\rm H}$ for simplification), $n_{\rm H^{+}}$ is the number density of neutral hydrogen atoms, and we assume that $n_{\rm H^{+}}=n_{\rm e}$ according to the conservation of charge. In addition, $m_{\rm e}$, $h$, and $\chi_{0}=13.6\,\rm{eV}$ are the electron mass, Planck constant, and ionization energy of the hydrogen atom, respectively.
Figure 1: A schematic diagram of the binary system in Section 3 (not to scale). Note that the darker regions in the disk have higher density. The magnetar continuously emits FRBs and orbits the Be star along an elliptical orbit. FRBs emitted by the magnetar at different positions along its orbit propagate along the LoS and reach the observer. The precession of the Be star disk leads to varying RMs among different FRBs.
3.2 Disk precession
In a Be star-NS binary, the supernova natal kick at the birth of the NS may lead to the misalignment of the orbital plane and the Be star disk plane (Brandt_1995MNRAS.274..461B),as well as a binary orbit with high eccentricity (van_den_Heuvel_1997ApJ...483..399V).
Following the method of Chen_2024ApJ...973..162C, we set a Cartesian coordinate system where the origin is at the center of the star, the $x$-axis points to the periastron of the binary star orbit, and the $z$-axis is perpendicular to the orbital plane (see Figure 1). The unit vectors of the $x$-axis, $y$-axis and $z$-axis, are denoted by $\vec{e}_{x}$, $\vec{e}_{y}$ and $\vec{e}_{z}$, respectively. The unit direction vector of the magnetar $\vec{e}_{\rm{mag}}$, and the angle between it and $x$ axis is $\psi$. The relation between the true anomaly $\psi$ and the time $t$ (orbital phase) is described by Kepler’s equation:
$\frac{2\pi t}{P_{\rm{orb}}}=2\rm{arctan}(\sqrt{\frac{1-\varepsilon}{1+\varepsilon}}\rm{tan}(-\psi/2))+\varepsilon\frac{\rm{sin}\psi\sqrt{1-\varepsilon^{2}}}{1+\varepsilon\rm{cos}\psi},$ (8)
where $\varepsilon$ is the eccentricity of the magnetar orbit.
The unit vector pointing from the Be star to the center of the magnetar can be expressed using the rotation matrix as follows:
$\vec{e}_{\rm{mag}}=\textbf{R}[\vec{e}_{z},-\psi]\vec{e}_{x}=\begin{bmatrix}\rm{cos\psi}\\ \rm{-sin\psi}\\ 0\end{bmatrix}$. (9)
The Be star disk precesses around a precession axis, which can be defined by a unit vector $\vec{e}_{\rm{pre}}$ and forms an angle $\varphi$ with the $z$ axis. To simplify the scenario, $\varphi$ is constrained in the $x$-$z$ plane. $\theta$ is the angle between $\vec{e}_{\rm{disk}}$ and $\vec{e}_{\rm{pre}}$. If the period of precession is $P_{\rm{pre}}$, $\vec{e}_{\rm{pre}}$ and $\vec{e}_{\rm{disk}}$ can be expressed as follows:
$\vec{e}_{\rm{pre}}=\textbf{R}[\vec{e}_{y},-\varphi]\vec{e}_{z}=\begin{bmatrix}\rm{-sin\varphi}\\ 0\\ \rm{cos\varphi}\end{bmatrix}$, (10)
$\vec{e}_{\rm{disk}}=\textbf{R}[\vec{e}_{\rm{pre}},\omega(t-t_{0})]\vec{e}_{\rm{disk,0}},$ (11)
where $\omega=2\pi/P_{\rm{pre}}$ is the angular velocity of precession, and $\vec{e}_{\rm{disk,0}}$ is the unit vector normal to the disk at the initial time $t_{0}$. To simplify, we define $\vec{e}_{\rm{disk,0}}$ as follows:
$\vec{e}_{\rm{disk,0}}=\textbf{R}[\vec{e}_{y},-(\frac{\pi}{2}+\theta+\varphi)]\vec{e}_{x}=\begin{bmatrix}\rm{-sin(\theta+\varphi)}\\ 0\\ \rm{cos(\theta+\varphi)}\end{bmatrix}$, (12)
where $\theta$ is the angle between $\vec{e}_{\rm{disk,0}}$ and $\vec{e}_{\rm{pre}}$.
In this coordinate system, the unit direction vector of LoS can be denoted as
$\vec{e}_{\rm{LoS}}=\begin{bmatrix}\rm{cos\lambda_{1}sin\lambda_{2}}\\ \rm{sin\lambda_{1}sin\lambda_{2}}\\ \rm{cos\lambda_{2}}\end{bmatrix}$, (13)
where the $\lambda_{1}$ or $\lambda_{2}$ are the angles between $\vec{e}_{\rm{LoS}}$ and $\vec{e}_{x}$ or $\vec{e}_{z}$.
4 Application to FRB 20201124A
In this section, we will use the precession disk model of the B star-magnetar binary in Section 3.2 to explain the variations in the RM data of FRB 20201124A.
4.1 Model parameters
In our model described in Section 3, the parameters are as follows
  1. .
    Stellar parameters: $_{\rm{NS}}, MM_{\star}$, and $R_{\star}$;
  2. .
[…]
  3. .
    Parameters of the binary orbit: $\varepsilon$ and $P_{\rm{orb}}$;
  4. .
    Parameters of the disk precession: $\varphi$ of $\vec{e}_{\rm{pre}}$, $\theta$ of $\vec{e}_{\rm{disk,0}}$, and $\omega$ of $\vec{e}_{\rm{disk}}$;
  5. .
    LoS: $\lambda_{1}$ and $\lambda_{2}$ of $\vec{e}_{\rm{LoS}}$.
For the Stellar parameters, their values are fixed in our calculations. For a B type star, its mass range is poorly constrained, spanning from a handful to a dozen of $M_{\odot}$. In this study, we fix $M\star$ as 8 $M_{\odot}$. Since the radius of B type stars ranges from a few to ten $R_{\odot}$, we adopt $R_{\star}=5\,R_{\odot}$. For the magnetar mass $M_{\rm{NS}}$, we use a typical NS mass of 1.6 $M_{\odot}$.
For the Disk properties, the isothermal temperature, the mean molecular weight, and the initial magnetic field in the disk are assumed to be $T=20,000\,\rm{K}$, $\mu=1$, and $\rm{B_{0}}=10\,\rm{G}$, respectively, referring to Wang_2022NatCo..13.4382W. Since the fitting is sensitive to the density profile of the disk, we set $\rho_{0}$ and $\beta$ as free parameters to achieve a better fitting result.
For the orbital parameters, i.e., $P_{\rm{orb}}$ and $\varepsilon$, we fix $P_{\rm{orb}}$ but treat $\varepsilon$ as a free parameter. According to the RM data of FRB 20201124A and our model, we anticipate that the maximum dips of E1 ($\sim$ MJD 59337) and E3 ($\sim$ MJD 59627) as seen in Figure 2 are very likely in the same phase of the magnetar orbital motion. Hence, our model expects that the time span between the two dips is an integer multiple of the orbital period $P_{\rm{orb}}$ and estimates the period to be $P_{\rm{orb}}=73$ days. With $P_{\rm{orb}}=73$ days, 4 orbital periods have been covered by the current FAST observations (see Figure 2). Since the orbital eccentricity $\varepsilon$ can significantly affect the profile of the RM variations, it is treated as a free parameter.
For the parameters of disk precession and LoS, since they are directly related to precession and RM integration, all of them are set as free parameters. In conclusion, our fitting parameters are $\varepsilon$, $\rho_{0}$, $\beta$, $\varphi$, $\lambda_{1}$, $\lambda_{2}$, $\theta$, and $\omega$.
4.2 Calculation methods and results
To compute the LoS integral of RM, we apply the method developed by Wang_2022NatCo..13.4382W: At each orbital phase, we take 1000 points along the LoS starting from the magnetar’s position, with a distance of $1\,R_{\odot}$ between two adjacent points. For each calculation point, the local plasma density and the magnetic field are calculated from Equations 1 and 4, respectively, and the expected $\Delta\rm{RM}$ is calculated from Equation 6.
Figure 2: The RM variation of FRB 20201124A in three FAST observation episodes, i.e., $\Delta\rm{RM}_{E1}$ (blue dots), $\Delta\rm{RM}_{E2}$ (maroon dots) and $\Delta\rm{RM}_{E3}$ (dark green dots). The ranges of observed RM of E1, E2 and E3 are shown as the blue, maroon, and green shadow regions, respectively. The fitting of the model presented in Section 3 is shown as the red curve, while the parameters of the fitting is shown in Table 1.
The orbital eccentricity $\varepsilon$ mainly affects the ratio of the peak and dip durations, as well as the amplitude of the $\Delta\rm{RM}$ curve. The larger $\varepsilon$, the greater the proportion of the peak in one period, and the more drastic the variation in the $\Delta\rm{RM}$ curve. The base mass density $\rho_{0}$ and the density radial profile $\beta$ can also influence the variation amplitude of the $\Delta\rm{RM}$ curve: increasing with increasing $\rho_{0}$ or decreasing $\beta$. For $\theta$, larger values lead to an increase in both the duration of the dip and its amplitude. $\varphi$ variations induce changes in the shape of the $\Delta\rm{RM}$ curve; positive $\varphi$ leads to a taller and wider peak and a tiny dip, while a more negative $\varphi$ results in a sharper dip (positive and negative $\varphi$ correspond to the positive and negative projections of $\vec{e}_{\rm{pre}}$ onto the x-axis in Figure 1, respectively). For $\omega$, it regulates how quickly the overall amplitude of the $\Delta\rm{RM}$ curve oscillates over each orbital period. The larger $\omega$, the faster the oscillation. For the LoS, variations in $\lambda_{1}$ significantly affect the proportion and amplitude of peaks and dips within one orbital period. A smaller $\lambda_{1}$ can result in a wider and deeper dip as well as a narrower and lower peak. For $\lambda_{2}$, its small perturbations can lead to significant changes in the amplitude of the $\Delta\rm{RM}$ curve. As $\lambda_{2}$ decreases, the amplitude of the $\Delta\rm{RM}$ curve increases — with the peak part increasing far more significantly than the dip part within an orbital period.
We fit the $\Delta\rm{RM}_{\rm{E1}}$, $\Delta\rm{RM}_{\rm{E2}}$ and $\Delta\rm{RM}_{\rm{E3}}$ data with our model. We obtained a set of parameters that are reasonable and well fit the RM variation data: $\varepsilon=0.69$, $\rho_{0}=3.49\times 10^{-15}\,\rm{g\,cm^{-3}}$, $\beta=4.09$, $\theta=0.07\,\rm{rad}$ ($\sim 4^{\circ}$), $\varphi=-0.7\,\rm{rad}$, $\omega=8.01\times 10^{-3}\,\rm{rad\,day^{-1}}$ (corresponding to a precession period of 784.71 days), $\lambda_{1}=-0.55\,\rm{rad}$, and $\lambda_{2}-\pi=-9.64\times 10^{-2}\,\rm{rad}$ (since $\lambda_{2}$ is very close to $\pi$, we adopt $\lambda_{2}-\pi$ to represent its magnitude). The results are presented in Table 1, and the fitting curve plotted using the best-fit parameters is shown in Figure 1 as the red curve.
Table 1: Fitting parameters of FRB 20201124A.
Parameter Best fit value
$M_{\star}\,(M_{\odot})$ 8
$R_{\star}\,(R_{\odot})$ 5
$M_{\rm{NS}}\,(M_{\odot})$ 1.6
$P_{\rm{orb}}\,\rm{(day)}$ 73
$\varepsilon$ 0.69
$\rho_{0}\,\rm{(g\,cm^{-3})}$ $3.49\times 10^{-15}$
$\beta$ 4.08
$\varphi\,\rm{(rad)}$ 0.7-0.7
$\theta\,\rm{(rad)}$ $0.07$
$\omega\,\rm{(rad\,day^{-1})}$ $8.01\times 10^{-3}$
$\lambda_{1}\,\rm{(rad)}$ 0.55-0.55
$\lambda_{2}-\pi\,\rm{(rad)}$ $9.64\times 10^{-2}$
5 Discussions and Conclusions
5.1 Discussions
Besides FRB 20201124A, notable RM variations have also been observed in several other FRB sources.
Unlike FRB 20201124A, these FRB sources do not have enough data from observations, making the constraint on disk precession loose. We apply our model to three of these FRBs, i.e. FRB 20180301A, FRB 20190520B, and FRB 20191106C, and also offer a prediction for their missing part of RM observations. The fittings and predictions of these sources are shown in Figure 3.
For each source, we present a set of parameters to plot two RM curves: the curve without precession and the curve of $P_{\rm{prec}}=10\times P_{\rm{orb}}$ (the two curves only differ in $\omega$ and $\theta$, since $\omega$ and $\theta$ are set to 0 for the curve without precession). Table 2 lists the parameter sets used for each source.
(a) FRB 20180301A
(b) FRB 20190520B
(c) FRB 20191106C
Figure 3: Three FRB sources with notable RM variation, each with model fitting RM curves of $P_{\rm{prec}}=10\times P_{\rm{orb}}$ (blue curve) and without precession (dotted red curve). The two curves in each subplot share a same set of parameters except $\theta$ and $\omega$, since these two parameters are both 0 for the no-precession case. Evolution of this two curves are shown for five orbital periods in each subplot. The observed $\Delta\rm{RM}$s is shown as the gray dots. The parameters of each FRB source are shown in Table 2.
In this work, we assume an ideal uniform precession of the Be star with a rigid decretion disk. The case would be, However, more complex in practice. In fact, whether a disk can behave as a rigid body depends on whether the interactions between its different parts are sufficiently strong — such as bending waves, viscous stresses, and self-gravity (Papaloizou_&Pringle_1983MNRAS.202.1181P; Papaloizou_&_Lin_1995ApJ...438..841P; Papaloizou_&_Terquem_1995MNRAS.274..987P; Ogilvie_1999MNRAS.304..557O; Ogilvie_2006MNRAS.365..977O; Bate_2000MNRAS.317..773B; Lubow_&_Ogilvie_2000ApJ...538..326L; Batygin_2012Natur.491..418B; Tremaine_2014MNRAS.441.1408T). Furthermore, according to Lai_2014MNRAS.440.3532L, the circumstellar disk of a massive star in a binary system might undergo precessional evolution, which is driven by the gravitational torques exerted by the companion star tidally. Larger stellar obliquity, i.e., the misalignment between the equatorial plane of the star and the orbital plane, might be induced by such evolution of the disk precession. However, such disk evolution is expected to be significant over a timescale of mega-years, so we do not consider it in our model.
According to the calculation of a rigid disk precession induced by a companion star in Lai_2014MNRAS.440.3532L, we estimate the precession period using the binary and disk parameters in Section 4.2. The disk with the density profile in Equation 1 can be regarded as a thin disk that spans from the inner radius $R_{\rm{in}}=R_{\star}$ to $R_{\rm{out}}=\eta R_{\star}$ with surface density:
$\Sigma(r)=\Sigma_{0}(r/R_{\star})^{-\beta+3/2},$ (14)
where $\Sigma_{0}$ is the disk surface density at $R_{in}$. The total disk mass is
$M_{d}=\int dM_{d}=2\pi\int_{R_{\rm{in}}}^{R_{\rm{out}}}r\Sigma(r)dr=\frac{4\pi\Sigma_{0}R_{\star}^{4}}{7-2\beta}(\eta^{7/2}-1),$ (15)
and for the disk angular momentum vector $\vec{L_{d}}=L_{d}\,\vec{e_{L_{d}}}$ (where $\vec{e_{L_{d}}}$ is the unit vector), can be calculated as
$L_{d}=\int rM_{d}(r)v_{k}(r)dr=\frac{2\pi\Sigma_{0}R_{\star}^{5/2}\sqrt{GM_{\star}}}{4-\beta}(\eta^{4-\beta}-1).$ (16)
The gravitational torque exerted on the disk by the external magnetar companion is:
$\vec{T_{d}}$ $=-\int^{R_{\rm{out}}}_{R_{\rm{in}}}dM_{d}\frac{3GM_{\rm{NS}}}{4a_{b}^{3}}\cos\theta\,\vec{e_{L_{b}}}\times\vec{e_{L_{d}}}$ (17)
$=\frac{3\pi GM_{\rm{NS}}\Sigma_{0}R_{\star}^{4}\cos\theta}{(11-2\beta)a_{b}^{3}}(\eta^{11/2-\beta}-1)\,\vec{e_{L_{b}}}\times\vec{e_{L_{d}}}$
$=\omega\vec{e_{L_{b}}}\times\vec{L_{d}},$
where $\vec{e_{L_{b}}}$ is the unit vector along the binary angular momentum axis, $a_{b}$ is the binary separation, $\theta$ is the inclination angle between $\vec{e_{L_{b}}}$ and $\vec{e_{L_{d}}}$, and $\omega$ is the angular velocity of the disk precession. Thus, the precession period of the disk, $P_{\rm{prec}}$, can be calculated as
$P_{\rm{prec}}$ $=2\pi/\omega=2\pi/|\vec{T_{d}}/\vec{L_{d}}|$ (18)
$=\frac{4\pi{a_{b}}^{3}}{3M_{\rm{NS}}R_{\star}^{3/2}cos\theta}\sqrt{\frac{M_{\star}}{G}}(\frac{11-2\beta}{4-\beta})(\frac{\eta^{4-\beta}-1}{\eta^{11/2-\beta}-1}).$
According to Section 4.2, $\beta=4$ is reasonable for the case of FRB 20201124A. When $\beta=4$, one has
$P_{\rm{prec}}^{\prime}$ $\simeq\frac{4\pi a_{b}^{3}}{M_{\rm{NS}}R_{\star}^{3/2}cos\theta}\sqrt{\frac{M_{\star}}{G}}\frac{ln(\eta)}{\eta^{3/2}-1}.$ (19)
$\eta$ shows an inverse correlation with the precession period $P_{\rm{prec}}^{\prime}$. We adopt the parameter values from Table 1, i.e., $\theta=0.07\,\rm{rad}$, $M_{\rm{NS}}=1.6M_{\odot}$, $M_{\star}=8M_{\odot}$, and $R_{\star}=5R_{\odot}$. We take the time averaged separation for $a_{b}$:
$\langle a_{b}\rangle_{t}\ =\frac{1}{P_{\rm{orb}}}\int_{0}^{P_{\rm{orb}}}a_{b}(t)dt=a(1+e^{2}/2),$ (20)
where $a$ and $P_{\rm{orb}}$ are the semi-major axes of the binary orbit. Here, we adopt $P_{\rm{orb}}=73\,\rm{day}$ and $e=0.69$ from Table 1 to calculate $a$ (see Equation 5) and $\langle a_{b}\rangle_{t}\$.
From the above formulae, we calculate that $\eta$ spans the range $[100,200]$, corresponding to $P_{\rm{prec}}^{\prime}$ in the range $[498,1220]$ days. The $\sim 900$-day period in Section 3 corresponds to $\eta\sim 141$. To estimate the impact of $\beta$ on the precession period, we fix $\eta$ to 125 in Equation 18. A $P_{\rm{prec}}$ in the range $[500,1200]$ days, corresponding to $\beta$ in the range $[3.71,4.21]$, and in this calculation $\beta=4$ is not included, as $\beta=4$ is a singularity in Equation 18 (an increase in $\beta$ leads to a corresponding increase in $P_{\rm{prec}}$). This is comparable with the $\beta$ in Table 1.
This means that using the model in Section 3 with reasonable $\eta$ and $\beta$, one can adequately explain the expected $\sim 785$-day precession period in Section 4.2.
5.2 Conclusions
In this work, we demonstrated that a Be star-magnetar binary system with a decretion stellar disk is suitable for explaining the peculiar RM variations observed by FAST from MJD 59307 to MJD 59658.
The RM in the first epoch exhibits a pattern of rising, a deeper drop, rising again, and then entering a steady phase. Such a pattern strongly supports that FRB 20201124A originates from a binary system consisting of a magnetar and a massive star with a decretion disk. In the later two observation epochs, the RM fluctuations decrease significantly compared with $\Delta\rm{RM_{E1}}$. The precession of the Be star along with its decretion disk is commonly used to explain the observed superorbital brightness oscillations of Be/X-ray binary system.
We propose that the long-term decrease in the observed RM variations can be interpreted by the superorbital modulation from a precessing stellar disk in a Be star-magnetar binary. By fitting the data, we obtainthe bestt parameters that fitthe RMM curves. The best parameters are presented in Section 4.
In addition, we use this model to explain a few other FRBs with similar dramatic RM variation, which can also be well explained and predicted. However, unlike FRB 20201124A, due to insufficient data sampling, the constraint of the model parameters remains limited.
With future continuous follow-up of these FRB sources with interesting RM variations, as well as the discovery of more new FRB sources with similar observational features, the understanding of the origin of such sources will be further advanced. These subsequent observations are also expected to enable better constraints on the model proposed in this paper.
We appreciate Dr. A. M. Chen for useful discussion of the precession. This work is supported by the National Natural Science Foundation of China under grants 12041306, 12473012, 12533005, 12173014 12494575, 12273009 and 12573044, and the National Key R&D Program of China (Nos. 2023YFC2205901, 2020YFC2201400, SQ2023YFC220007). W.H.Lei. acknowledges support from science research grants from the China Manned Space Project with NO.CMS-CSST-2021-B11.
Table 2: Parameters of fitting curves for the three FRB sources in Figure 3
Parameters FRB 20180301A FRB 20190520A FRB 20191106C
$M_{\star}\,(M_{\odot})$ 8 8 8
$R_{\star}\,(R_{\odot})$ 5 5 5
$M_{\rm{NS}}\,(M_{\odot})$ 1.6 1.6 1.6
$\varepsilon$ 0.5 0.5 0.7
$P_{\rm{orb}}\,\rm{(day)}$ 2500 800 4500
$\rho_{0}\,\rm{(g\,cm^{-3})}$ $9.0\times 10^{-14}$ $6.2\times 10^{-12}$ $4.0\times 10^{-13}$
$\beta$ 2.5 3.0 3.0
$\varphi\,\rm{(rad)}$ -0.785 -0.785 -0.780
$\theta\,\rm{(rad)}$ 0.1 0.01 0.05
$\omega\,\rm{(rad\,day^{-1})}$ $1.26\times 10^{-3}$ $3.93\times 10^{-3}$ $6.98\times 10^{-4}$
$\lambda_{1}\,\rm{(rad)}$ -0.61 -0.40 -0.35
$\lambda_{2}\,\rm{(rad)}$ 3 2.98 2.6